# Linear Model of Aggregated Homogeneous Energy Storage Elements with Realizable Dispatch Guarantees

### v4

@@ Abstract
To optimize the dispatch of batteries, a model is required that can predict the state of charge (SOC) trajectory for a chosen open-loop power schedule to ensure admissibility (i.e., that schedule can be realized). However, battery dispatch optimization is inherently challenging since batteries cannot simultaneously charge and discharge, which begets a non-convex complementarity constraint. In this paper, we develop a novel composition of energy storage elements that can charge or discharge independently and provide a sufficient linear energy storage model of the composite battery. This permits convex optimization of the composite battery SOC trajectory while guaranteeing admissibility of the resulting (aggregated) power schedule and its disaggregation to the individual energy storage elements.
@@
Battery, energy storage, optimization, convex, priority-based control policy, complementarity conditions.
@@ I Introduction
Battery energy storage systems (BESS) are increasingly deployed to help facilitate the grid’s uptake of variable renewable energy and improve reliability  [Battery_review_paper, BESS_simultaneous_participation, BESS_renewables]. In many applications, BESS dispatch involves coordinating $N$ battery racks/modules that together comprise a single BESS unit (e.g., DC-coupled elements), $N$ BESS container units in a utility-scale battery storage facility (e.g., AC-coupled elements), or $N$ aggregated BESS units coordinated as a virtual battery/power plant (VPP). However, when optimizing BESS dispatch, the individual elements that make up the BESS are often lumped together into a single composite energy storage asset. The composite dispatch is then optimized and the resulting power schedule is shared equally among the $N$ battery elements. An optimal composite dispatch represents a non-convex problem since the complementarity constraint is formulated with a non-linear constraint or by the use of binary variables [MILP_for_CC, prat_2024].
To obtain a convex relaxation of the dispatch problem, the complementarity constraint is typically omitted [prat_2024]. The resulting power trajectories can then exhibit simultaneous charging and discharging, which cannot be implemented. To recover near-term feasibility, only the net power trajectory is applied. As a result, the state of charge (SOC) becomes a provable lower bound on the SOC when the schedule is realized [RobustNawaf, almassalkhi_model-predictive_2015]. This can lead to unexpected SOC saturation, resulting in an un-realizable power trajectory. Extensive research has investigated the conditions under which the convex relaxation’s state of charge (SOC) coincides with the actual realized SOC. Authors in [K.Baker_DCOPF_No_CC] avoid the use of the complementarity constraint when applied to DCOPF but add a penalty term to the objective function to ensure the exactness of the relaxed model. Ex-post sufficient conditions under which the relaxation becomes exact for an economic dispatch problem are provided in [Sufficient_Cond_for_Relax, Extended_Sufficient_Cond_for_Relax]. Ex-ante conditions for exactness of the relaxed problem are presented in [Lin_Exact_Cond_Ex_Ante], assuming that certain constraints on battery power limits, as well as initial and maximum energy capacities, are met. Furthermore, [positive_prices, Jakob_Pos_Quadratic_Cost, prat_2024] shows that for economic dispatch problems with positive linear and quadratic cost functions, the convex relaxation at optimality would be exact with no simultaneous charging and discharging.
However, the exactness of the relaxed model does not hold in general which may lead to battery saturation and inadmissible power trajectories. A simple example of this is the case with negative prices as shown in [Baldick_Neg_Price_Bad]. For a single time step, this problem can be alleviated by developing tighter relaxations [D.Pozo_paper, elgersma2024], which can also heuristically be applied to time series, albeit without realizability guarantees.
Figure 1: Comparing different composite models after implementing an open-loop predicted dispatch schedule. The relaxed model dispatch is inadmissible, while the robust linear model is admissible but suboptimal. However, this paper’s model achieves near-optimal performance and is admissible.
Additionally, authors in  [RobustNawaf, Mazen_CDC_paper], an alternative robust linear battery model is presented which is a convex inner approximation of the non-convex formulation that sacrifices optimality for guaranteed admissibility. The robust linear model omits the complementarity constraint but envelopes the actual SOC trajectories to guarantee that any feasible power trajectory is realizable. However, the envelopes beget conservativeness as a function of efficiencies and prediction horizon, which limits practicality over longer horizons. In summary, previous work in energy storage optimization that circumvents the complementarity condition either has been application-specific and holds only if specific criteria on the objective function and/or the battery parameters are met, or presents convex restrictions that may yield significantly sub-optimal solutions. Furthermore, previous work has generally (and often, implicitly) assumed an equal power-sharing policy, which limits how individual elements can respond under dynamic conditions.
Thus, this manuscript revisits the modeling of BESS aggregations by presenting a novel composite model $N$ identical and independently controllable battery elements, where simultaneous charging and discharging power trajectories at the composite level can be implemented by inhomogeneous disaggregation to the individual elements. The main contributions of the paper are as follows:
@@
    We present a novel linear model that provides a sufficient representation of aggregations of homogeneous battery storage elements.
@@
    We develop a priority-based strategy, based on element SOC, to perform the composite-to-individual disaggregation.
@@
    The proposed model guarantees that all individual element-wise power limits, energy limits, and complementarity constraints will be satisfied after disaggregation.
Figure 1 illustrates the open-loop implementation of the dispatch schedules obtained from the proposed model, the relaxed model, and the robust linear model presented in [RobustNawaf] when used to optimize composite battery dispatch to track a given power reference signal. The relaxed model dispatch causes the battery to saturate leading to poor tracking performance. The robust linear model dispatch is realizable; however, yields sub-optimal tracking performance. The proposed model yields near-optimal results and guarantees realizable power trajectories under an unequal power-sharing policy.
The rest of the paper is organized as follows. Section II presents the composite battery model and the associated constraints. Section III presents the proposed disaggregation strategy and the sufficient conditions for the admissibility of the composite power trajectory. Section IV presents simulation results to demonstrate the performance of the proposed model and disaggregation strategy. Finally, Section V concludes the paper.
@@ II Composite Battery Model
In this section, we will present element-wise and composite models of BESS systems. We consider BESS assets composed of an aggregated generic energy storage asset. An example of such a system is shown in Fig 2. It is composed of $N$ identical elements.
For element $i\in\{1,\ldots,N\}$, the charge/discharge power is set in controller time steps $l$ with duration $\delta t$. The charge/discharge power between $l$ and $l+1$, the charge/discharge efficiency and the SOC at timestep $l$ are denoted by $P^{i}_{\text{c/d}}(l)$, $\eta_{\text{c/d}}$, and $E^{i}(l)$, respectively. A realizable BESS dispatch must satisfy power limits (${P}_{\text{c/d,max}}$), energy limits (${E}_{\text{max}}$), and complementarity constraints for all battery elements ($\forall i=1,\ldots,N$) and time steps ($\forall l=0,\ldots,L-1$). Note that all elements have identical power and energy limits. This results in (1), where $E^{i}_{0}$ is the initial SOC of battery element $i$:
@@
$0=P_{\text{c}}^{i}(l)P_{\text{d}}^{i}(l).$ (1e)
The scheduling of the composite battery is done with (optionally) longer scheduler time steps of duration $\Delta t=M\delta t,M\in\mathbb{N}^{+}$. We denote these scheduling time steps with index $[k]$ in square brackets, where $k=0,\ldots,K-1$ and $L=MK$. Constant power in step $k$ implies the following constraint for all $l,k$ with $k=\lfloor\frac{l}{M}\rfloor$:
@@
Typically, it is assumed that a composite BESS system is modeled and operated under balanced conditions: the dispatcher (e.g., EMS) maintains equal SOC ($E_{0}^{i}=E_{0}~{}~{}\forall i$) and adopts an equal power sharing policy across its elements. This results in a composite battery with power and energy limits of $N{P}_{\text{c/d,max}}$ and $N{E}_{\text{max}}$ as shown in (3):
@@
Under the equal power sharing policy, a single complementarity constraint (3e) at the composite battery level is sufficient to satisfy (1e) for all $i$ and all $l$. Under balanced operation, a feasible composite dispatch based on (3) satisfies all elements’ constraints in (1).
@@
where the cutting plane in (4b) has been added to tighten the relaxation of the problem. However, such relaxations can result in an optimal power schedule with simultaneous charging and discharging, which cannot be implemented under equal load sharing. Tighter relaxations have been developed [D.Pozo_paper, elgersma2024] to reduce the prevalence of such solutions in real-world problems.
Nevertheless, there are cases where simultaneous charging and discharging (which effectively ‘discards’ energy if $\eta_{c}\eta_{d}<1$) is a desirable outcome of a battery dispatch schedule. For example, this may be the case when energy prices are negative, when the storage unit tracks a dispatch signal, or as byproduct of minimum up- or downtime constraints in a centrally dispatched system.
@@ III-A Priority Stack Controller (PSC)
For every controller time-step $l$, the priority stack controller (PSC) sorts all $N$ elements based on the SOC. The element with the lowest SOC has the highest charging priority while the element with the highest SOC has the highest discharging priority. When selecting the highest-priority element for charging/discharging that element’s power is maximized, i.e., $P^{i}_{\text{c/d}}(l)=P_{\text{c/d,max}}$, before actuating another charging/discharging element from the stack. This process is repeated every time-step $l$, adapting to changing element SOCs. The priority stack results in $N_{\text{c}}$ elements with lowest SOCs charging and $N_{\text{d}}$ elements with highest SOCs discharging, as follows:
@@
The PSC is summarized in Algorithm 1 and begets Property 1.
@@
6:    Update SOC for all elements: $E^{i}(l)\rightarrow E^{i}(l+1)$
@@ III-B Guaranteeing Element Complementarity Constraints
Define $\Gamma_{\text{c}}:=\lfloor\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}\rfloor$ and $\Delta_{\text{c}}:=\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}-\Gamma_{\text{c}}\in[0,1)$. Then $\Big{\lceil}\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}\Big{\rceil}\leq\Gamma_{c}+1$ and, from (7), we get
@@
Figure 3: Feasible region for $N=10$ and $P_{\text{c/d,max}}=0.5$MW (left) and $N=30$ and $P_{\text{c/d,max}}=0.17$MW (right). SOC limits are omitted here.
III-C Relating Element and Composite Energy Limit Constraints
Similarly, we want to relate element energy limits to those in the composite formulation. Under PSC, the battery elements are charged/discharged individually every $l$, which results in differing SOC trajectories. The spread of element SOCs can impact realizability of the composite dispatch when individual elements hit their SOC limits. Constraint violations can be avoided by quantifying the maximum possible SOC spread under PSC and using that as an energy buffer applied to the composite battery energy limits.
@@ III-C Relating Element and Composite Energy Limit Constraints
Define the maximum difference in SOC across all elements as $\Delta E(l):=\max_{i,j}\{|E^{i}(l)-E^{j}(l)|\}$ (i.e., the SOC spread). Consider element SOCs $E^{i}(l)$ and $E^{j}(l)$ and, without loss of generality, let $E^{i}(l)\geq E^{j}(l)$. Using (1a) the following is obtained,
@@
where the inequality in (12b) uses $E^{i}(l)-E^{j}(l)\geq 0$. Considering all $i,j$, we can bound $\Delta E(l+1)\leq\max\left\{\Delta E(l),\varepsilon\right\}$. Moreover, we see that $\Delta E(l)\leq\varepsilon\Rightarrow\Delta E(l+1)\leq\varepsilon$, so it follows that for a composite battery with initial SOC spread $\Delta E(0)\leq\varepsilon$ (condition (i)), we maintain $\Delta E(l)\leq\varepsilon,\forall l$.
Because energy linearly increases/decreases within each scheduling time step $k$ under PSC, (9) implies its equivalent for controller time steps $l$: $N\varepsilon\leq E(l)\leq N(E_{\text{max}}-\varepsilon)$. Dividing by $N$ begets $\varepsilon\leq\overline{E}(l)\leq E_{\text{max}}-\varepsilon$, where $\overline{E}(l):=\sum_{j}E^{j}(l)/N$ is the average element SOC. Because the SOC spread $\Delta E(l)\leq\varepsilon$, we must have $\overline{E}(l)-\varepsilon\leq E^{i}(l)\leq\overline{E}(l)+\varepsilon,\forall i$. Combined, these inequalities show that (1c) holds for all $i$ and all $l$.

### v6

# Linear Aggregate Model for Realizable Dispatch of Homogeneous Energy Storage

@@ Abstract
To optimize the dispatch of batteries, a model is required that can predict the state of energy (SOE) trajectory for a chosen open-loop power schedule to ensure admissibility (i.e., that schedule can be realized). However, battery dispatch optimization is inherently challenging when batteries cannot simultaneously charge and discharge, which begets a non-convex complementarity constraint. In this paper, we develop a novel composition of energy storage elements that can charge or discharge independently and provide a sufficient linear energy storage model of the composite battery. This permits convex optimization of the composite battery dispatch while ensuring the admissibility of the resulting (aggregated) power schedule and its disaggregation to the individual elements.
@@
Battery, energy storage, convex optimization, priority-based control, complementarity constraints.
@@ I Introduction
Battery energy storage systems (BESS) are increasingly deployed to help facilitate the grid’s uptake of variable renewable energy and improve reliability  [Battery_review_paper, BESS_simultaneous_participation, BESS_renewables]. In many applications, BESS dispatch involves coordinating large numbers of energy storage elements. Examples include battery racks that comprise a single BESS asset, BESS container assets in a utility-scale battery storage facility as shown in Fig 1, or aggregated BESS units coordinated as a virtual power plant (VPP).
Optimally scheduling the charging and discharging of storage elements requires solving a non-convex optimization problem. Because most storage technologies do not permit simultaneous charging and discharging, complementarity constraints are usually required to force the product of the charging and discharging powers to be equal to zero at each time step, and for each element. This nonlinearity is often implemented using binary variables [MILP_for_CC, prat_2024]. The non-convexity of the composite BESS dispatch can present a computational challenge when incorporating BESSs into more complex, large-scale power system optimization problems, such as optimal power flow (OPF) and economic dispatch [qi2025locationalenergystoragebid].
This motivates the need for a convex BESS model. A common approach is to obtain a convex relaxation by omitting the complementarity constraint, which results in a predicted state of energy (SOE) that is a provable lower bound on the realized SOE and can lead to unexpected SOE saturation and an inadmissible dispatch [RobustNawaf, almassalkhi_model-predictive_2015]. Additionally, the relaxation permits solutions that simultaneously charge and discharge [prat_2024].
At first glance, such solutions may appear irrational because charging/discharging inefficiencies result in unnecessary dissipation of energy. Nevertheless, there are cases where such solutions are a rational outcome of a battery dispatch schedule, e.g., when energy prices are negative [Baldick_Neg_Price_Bad], or as a byproduct of minimum up- or downtime constraints in a centrally dispatched system. Any dissipative effects on battery aging can be captured by adding a cycling cost to the objective function, so that the benefit must outweigh the cost. Sufficient conditions under which simultaneously charging and discharging is optimal (i.e., rational) are given in [prat_2024].
Conversely, there are cases where simultaneous charging and discharging is guaranteed not to occur. Sufficient ex-post conditions for an exact relaxation in economic dispatch are provided in [Sufficient_Cond_for_Relax]. Ex-ante conditions for exactness of the relaxed problem are presented in [Lin_Exact_Cond_Ex_Ante], assuming that battery power limits, as well as initial and maximum energy capacities, satisfy certain conditions. Furthermore, [positive_prices, Jakob_Pos_Quadratic_Cost, prat_2024] show that for specific economic dispatch formulations with positive linear or quadratic cost functions, the relaxation at optimality would be exact with no simultaneous charging and discharging.
Various alternative problem formulations exist to reduce the reliance on the computationally demanding complementarity constraint. For a single time step, tighter relaxations have been proposed [D.Pozo_paper, elgersma2024], which can also heuristically be applied to time series, albeit without realizability guarantees. Authors in [K.Baker_DCOPF_No_CC] avoid the use of the complementarity constraint when applied to DC Optimal Power Flow (DCOPF) but add a penalty term to the objective function to ensure the exactness of the relaxed model. Instead of modifying objective, the authors in  [RobustNawaf, Mazen_CDC_paper] present an alternative robust linear battery model which represents a convex restriction of the non-convex formulation that sacrifices optimality for guaranteed admissibility. The robust linear model omits the complementarity constraint but envelopes the actual SOE trajectories to guarantee that any feasible power trajectory is realizable. However, the envelopes beget conservativeness as a function of energy conversion efficiencies and the prediction horizon, which limits practicality.Now, a composite battery made up of identical storage elements can be scheduled with various methods. One way is to model and optimize the schedule of each storage element individually, requiring many complementarity constraints [equal_unequal]. A simpler and computationally efficient approach is to use a static equal power sharing policy where all elements charge or discharge equally [equal_unequal, equal_unequal2]. The elements can then be aggregated into a composite battery whose dispatch is optimized, but still requires one non-convex complementarity constraint. In summary, previous work in energy storage optimization that circumvents the complementarity condition either has been application-specific and holds only if specific criteria on the objective function and/or the battery parameters are met, or presents convex restrictions that may yield significantly sub-optimal solutions. Furthermore, previous work has generally (and often, implicitly) assumed an equal power-sharing policy, which limits how elements can be dispatched. This manuscript revisits the modeling of aggregations of storage units by presenting a novel composite model of $N$ identical and independently controllable battery elements, where simultaneous charging and discharging power trajectories at the composite level can be implemented by inhomogeneous disaggregation to the individual elements. Throughout, BESS aggregations are used as an example, but the proposed work can be extended to different energy storage assets. The main contributions of the paper are as follows:
@@
    We present a novel linear model that provides a sufficient representation of aggregations of homogeneous energy storage elements. For large numbers of elements or and small control time steps, the representation asymptotically approaches the true feasibility envelope.
@@
    We develop a priority-based strategy, based on element SOE, to perform composite-to-elements disaggregation.
@@
    We prove that all solutions of the linear aggregate model are realizable using the priority-based strategy, respecting element-wise power limits, energy limits, and complementarity constraints after disaggregation.
Next, we develop the composite battery model in Section II. Section III presents the proposed disaggregation strategy and sufficient conditions for the admissibility of a composite power trajectory. Simulation results that demonstrate the performance of the proposed model and disaggregation strategy make up Section IV. Conclusion and future work are in Section V.
@@ II Composite Battery Model
In this section, we will present element-wise and composite models of BESS systems, composed of $N$ identical and independently controllable elements (e.g., as shown in Fig 1). For element $i\in\{1,\ldots,N\}$, the charge/discharge power is set in controller time steps $l$ with duration $\delta t$. The charge/discharge power between $l$ and $l+1$, the charge/discharge efficiency and the SOE at timestep $l$ are denoted by $P^{i}_{\text{c/d}}(l)$, $\eta_{\text{c/d}}$, and $E^{i}(l)$, respectively. A realizable BESS dispatch must satisfy power limits (${P}_{\text{c/d,max}}$), energy limits (${E}_{\text{max}}$), and complementarity constraints for all battery elements ($\forall i=1,\ldots,N$) and time steps ($\forall l=0,\ldots,L-1$). Note that all elements have identical power and energy limits. This results in (1), where $E^{i}_{0}$ is the initial SOE of battery element $i$:
@@
$0=P_{\text{c}}^{i}(l)P_{\text{d}}^{i}(l),$ (1e)
where (1c)-(1e) represent energy limits, power limits and, complementarity constraints, respectively. The scheduling of the composite battery is done with (optionally) longer scheduler time steps of duration $\Delta t=M\delta t,M\in\mathbb{N}^{+}$. We denote these scheduling time steps with index $[k]$ in square brackets, where $k=0,\ldots,K-1$ and $L=MK$. Constant power in step $k$ implies the following constraint for all $l,k$ with $k=\lfloor\frac{l}{M}\rfloor$:
@@
Typically, it is assumed that a composite BESS system is modeled and operated under balanced conditions: the dispatcher (e.g., EMS) maintains equal SOE ($E_{0}^{i}=E_{0}~{}~{}\forall i$) and adopts an equal power sharing policy across its elements. This results in a composite battery with power and energy limits of $N{P}_{\text{c/d,max}}$ and $N{E}_{\text{max}}$ as shown in (3):
@@
Under equal power sharing, a complementarity constraint (3e) at the composite battery level is sufficient to satisfy (1e) for all $i$ and all $l$. Under balanced operation, a composite dispatch satisfying (3) would satisfy (1) for all $i$ and all $l$..
@@
where the cutting plane in (4b) is one of the facets of the convex hull of the feasible set defined by (3), typically used to tighten the relaxation. Even tighter relaxations have been developed [D.Pozo_paper, elgersma2024] to reduce the prevalence of solutions with simultaneous charging and discharging in real-world problems. However, such relaxations can result in an optimal power schedule with simultaneous charging and discharging, which cannot be implemented under equal load sharing.
@@ III-A Priority Stack Controller (PSC)
For every controller time-step $l$, the priority stack controller (PSC) sorts all $N$ elements based on the SOE. The element with the lowest SOE has the highest charging priority, while the element with the highest SOE has the highest discharging priority. Note that similar control schemes have been proposed in [karan_kalsi] and investigated in a real-time operational setting in [equal_unequal, equal_unequal2].
The priority stack results in $N_{\text{c}}$ elements with lowest SOEs charging and $N_{\text{d}}$ elements with highest SOEs discharging, as follows:
@@
Disaggregating the composite charging trajectory for $P_{\text{c}}[k]>0$, the PSC sets $P^{i}_{\text{c}}(l)=P_{\text{c,max}}$ for the $N_{\text{c}}(l)-1$ elements with the lowest SOE and the next-lowest element is assigned $P^{i}_{\text{c}}(l)=P_{\text{c}}[k]-(N_{\text{c}}(l)-1)P_{\text{c,max}}$. Discharging is implemented analogously. The PSC is summarized in Algorithm 1 and begets Property 1.
@@
6:    Update SOE for all elements: $E^{i}(l)\rightarrow E^{i}(l+1)$
@@ III-B Guaranteeing Element Complementarity Constraints
Define $\Gamma_{\text{c}}:=\lfloor\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}\rfloor$ and $\Delta_{\text{c}}:=\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}-\Gamma_{\text{c}}\in[0,1)$ to be the integer and fractional components of the quantity $\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}$. Then $\Big{\lceil}\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}\Big{\rceil}\leq\Gamma_{c}+1$ and, from (7), we get
@@
Figure 2: Feasible region for $N=10$ and $P_{\text{c/d,max}}=0.5$MW (left) and $N=30$ and $P_{\text{c/d,max}}=0.17$MW (right). SOE limits are omitted here.
III-C Relating Element and Composite Energy Constraints
Similarly, we want to relate element energy limits to those in the composite formulation. Under PSC, the battery elements are charged/discharged individually every $l$, which results in differing SOE trajectories. The spread of element SOEs can impact the realizability of the composite dispatch when individual elements hit their SOE limits. Constraint violations can be avoided by quantifying the maximum possible SOE spread under PSC and using that as an energy buffer applied to the composite battery energy limits.
@@ III-C Relating Element and Composite Energy Constraints
Define the maximum difference in SOE across all elements as $\Delta E(l):=\max_{i,j}\{|E^{i}(l)-E^{j}(l)|\}$ (i.e., the SOE spread). Consider element SOEs $E^{i}(l)$ and $E^{j}(l)$ and, without loss of generality, let $E^{i}(l)\geq E^{j}(l)$. Using (1a) the following is obtained,
@@
where the inequality in (12b) uses $E^{i}(l)-E^{j}(l)\geq 0$. Considering all $i,j$, we can bound $\Delta E(l+1)\leq\max\left\{\Delta E(l),\varepsilon\right\}$. Moreover, we see that $\Delta E(l)\leq\varepsilon\Rightarrow\Delta E(l+1)\leq\varepsilon$, so it follows that for a composite battery with initial SOE spread $\Delta E(0)\leq\varepsilon$ (condition (i)), we maintain $\Delta E(l)\leq\varepsilon,\forall l$.
Because energy linearly increases/decreases within each scheduling time step $k$ under PSC, (9) implies its equivalent for controller time steps $l$: $N\varepsilon\leq E(l)\leq N(E_{\text{max}}-\varepsilon)$. Dividing by $N$ begets $\varepsilon\leq\overline{E}(l)\leq E_{\text{max}}-\varepsilon$, where $\overline{E}(l):=\sum_{j}E^{j}(l)/N$ is the average element SOE. Because the SOE spread $\Delta E(l)\leq\varepsilon$, we must have $\overline{E}(l)-\varepsilon\leq E^{i}(l)\leq\overline{E}(l)+\varepsilon,\forall i$. Thus, (1c) holds for all $i$ and all $l$.